OpenStudy (anonymous):

write the slope-intercept form of the equation of the line described (0,5), perp. to y=-2/5x+2

sam (.sam.):

perpendicular, \[m_1m_2=-1\] \[-2/5 m_2=-1\] \[m_2=5/2\] \[y-y_1=m(x-x_1)\] y-5=5/2(x-0) y=5x/2 +5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I have one more

OpenStudy (anonymous):

write the slope-intercept form of the equation of the line through the given point with the given slope (-1,1) slope =-6

sam (.sam.):

y-y1=m(x-x1)

OpenStudy (anonymous):

wich is right above it.

sam (.sam.):

y-1=(-6)(x--1) y=-6x-6+1 y=-6x-5

OpenStudy (anonymous):

thanks so much

sam (.sam.):

First question answer: y=5x/2 +5 Second question answer: y=-6x-5